# MuRating: A High Quality Data Selecting Approach to Multilingual Large Language Model Pretraining

### v1

@@ A.1 Different Annotation Method
Askllm We adopt the approach from [43] and use the following prompt to query Flan-T5-xxl [8] for annotation 300,000 document pairs.
@@
A.2 Training details of MuRater and training accuracy
We adopt the XLM-RoBERTa architecture encoder model BGE-M3 [6] as the foundation of our multilingual rating model, MuRater, and fine-tune it by appending a linear regression head to the transformer output to predict quality scores. The fine-tuning process employs a confidence margin threshold of 50%, defined as $|p_{A}-p_{B}|=|2p_{B\succ A}-1|$ for a prediction between text pairs $(t_{A},t_{B})$ [55]. Fine-tuning is conducted over 3 epochs with a batch size of 512 and a learning rate of $2\times 10^{-5}$. We set $\lambda$ to 0.5. Performance on held-in and held-out sets is summarized in Table 3. Notably, BGE-M3 supports over 100 languages and leverages large-scale multilingual unsupervised data to learn a shared semantic space, making it particularly effective for multilingual and cross-lingual retrieval and rating tasks.
@@ A.2 Training details of MuRater and training accuracy
Appendix B Experiment Setup Details
B.1 Dataset
We apply the Fineweb and Fineweb-2 [37] as our training corpus. We employ NVIDIA’s multilingual domain classifier [^7] https://huggingface.co/nvidia/multilingual-domain-classifier to label the domain distribution of our dataset. The Figures 7-11 depict the domain distributions before and after applying MuRater-based selection. The results indicate that MuRater consistently prioritizes knowledge-intensive domains, such as People and Society, Health, and Science. These domains are generally characterized by well-organized content and high informational density, which are advantageous for the pretraining of LLMs. Nevertheless, the selected domain distributions are not same across languages, primarily due to significant domain differences inherent in their respective source corpora.
@@ B.1 Dataset
B.2 Baselines
We follow the same annotation procedure for the English datasets of QuRater, AskLLM, DCLM, and FineWeb-Edu as described in Appendix A. For QuRater-M, we apply the same prompting approach (also detailed in Appendix A) and instruct GPT-4o to annotate 300,000 multilingual pairs, focusing exclusively on content regardless of the language. We then fine-tune the multilingual QuRater baseline using both English and multilingual data, leveraging the BGE-M3 model [6] and the identical training hyperparameters outlined in Appendix A.
B.3 Model Architecture
We utilize a transformer architecture based on the LLaMA-2 model [50], configured to contain approximately 1.2 billion parameters. Models are randomly initialized before pretraining. The detailed information for the model configuration and training hyperparamters is shown in Table 4.
@@ B.3 Model Architecture
B.4 Benchmarks
Appendix C Evaluation Benchmarks
All task evaluations are conducted using the lm-evaluation-harness framework [14]. For English in-context learning tasks, we use the following benchmakrs:
  • ARC-Easy and ARC-Challenge [10] (25-shot): Multiple-choice science questions from grade school exams, assessing models’ ability to apply scientific knowledge and reasoning.
  • SciQ [54] (0-shot): Crowdsourced multiple-choice science questions covering physics, chemistry, and biology, designed to evaluate scientific understanding.
  • LogiQA [29] (0-shot): Logical reasoning questions derived from Chinese civil service exams, testing deductive reasoning capabilities.
  • TriviaQA [21] (5-shot): Reading comprehension dataset with question-answer pairs authored by trivia enthusiasts, accompanied by evidence documents.
  • BoolQ [9] (5-shot): Yes/no questions with associated passages, evaluating models’ ability to answer naturally occurring questions.
@@ Appendix C Evaluation Benchmarks
  • HellaSwag [61] (10-shot): Sentence completion tasks requiring commonsense inference to select the most plausible continuation.
  • PIQA [3] (5-shot): Physical commonsense reasoning questions, focusing on everyday tasks and interactions.
  • OpenBookQA [32] (10-shot): Multiple-choice questions based on elementary science facts, requiring both factual knowledge and reasoning.
  • WinoGrande [44] (5-shot): Pronoun resolution tasks designed to test commonsense reasoning at scale.
Additionally, two knowledge-intensive tasks are evaluated:
  • Natural Questions (NQ) [24] (5-shot): Real user questions paired with answers from Wikipedia, assessing open-domain question answering.
  • MMLU [19] (5-shot): A benchmark covering 57 subjects across various domains, measuring multitask language understanding.
For evaluating translated benchmarks, we use the MuBench dataset [18] and conduct evaluations across 18 languages present in our training set. In the multilingual setting, we evaluate:
@@
  • StoryCloze [33] (0-shot): Narrative understanding task where models choose the correct ending to a four-sentence story.
  • BMLAMA [41] (0-shot): Multilingual factual knowledge probing dataset, assessing cross-lingual consistency in language models.
  • XCOPA [40] (5-shot): Causal commonsense reasoning tasks translated into multiple languages, evaluating cross-lingual inference.
  • XNLI [11] (5-shot): Cross-lingual natural language inference benchmark, testing entailment and contradiction detection.
  • XWinograd [49] (5-shot): Multilingual pronoun resolution tasks, assessing commonsense reasoning across languages.
  • FLORES [15] (5-shot): Multilingual machine translation benchmark, evaluating translation quality across diverse languages.
@@
C.1 Translation
The translation prompts is
C.2 Human translation quality evaluation
To assess the translation quality of GPT-4o outputs, we engaged professional human translators to evaluate a subset of the generated translations. Each language translation was reviewed by a single expert. Evaluators were compensated at a rate of \$16 per hour, with each assessment session lasting approximately 4 hours. the annotation criteria for translation quality is shown below.
C.3 Pointwise Score
Appendix D Detailed Results
D.1 English Detailed Results
@@
D.2 Multilingual Detailed Results
We display the detailed results of each benchmark and each language below.
@@ D.2 Multilingual Detailed Results
Appendix E Case Study
@@
Figure 14: Sampled training examples of France with quality ratings at different score range

### v2

@@ 6.1 Different Annotation Method
AskLLM We adopt the approach from [sachdeva2024train] and use the following prompt to query Flan-T5-xxl [chung2024scaling] for annotation 300,000 document pairs.
@@
6.2 Training details of MuRater and training accuracy
We adopt the XLM-RoBERTa architecture encoder model BGE-M3 [bge-m3] as the foundation of our multilingual rating model, MuRater, and fine-tune it by appending a linear regression head to the transformer output to predict quality scores. The fine-tuning process employs a confidence margin threshold of 50%, defined as $\epsilon=|p_{A}-p_{B}|=|2p_{B\succ A}-1|$ for a prediction between text pairs $(t_{A},t_{B})$ [wettig2024qurating]. Fine-tuning is conducted over 3 epochs with a batch size of 512 and a learning rate of $2\times 10^{-5}$. We set $\lambda$ to 0.5. Performance on held-in and held-out sets is summarized in Table 3. Notably, BGE-M3 supports over 100 languages and leverages large-scale multilingual unsupervised data to learn a shared semantic space, making it particularly effective for multilingual and cross-lingual retrieval and rating tasks.
@@ 6.2 Training details of MuRater and training accuracy
6.3 Translation
The translation prompts is
6.4 Human translation quality evaluation
To evaluate the translation quality of GPT-4o outputs, we employed professional human translators to assess a selected subset of the generated texts. All evaluators possessed CEFR C1-level or higher proficiency in both English and the respective target language. Each language translation was reviewed by a single expert. Evaluators were compensated at a rate of \$16 per hour, with each assessment session lasting approximately 4 hours. the annotation criteria for translation quality is shown below.
6.5 Translation Quality Asseement
Figure 6: Translation average scores of various languages.
We assess the quality of our translations through human evaluation. Expert annotators are provided with 50 pairs of source and translated texts and asked to rate translation quality on a scale mentioned above. As shown in Figure 6, the overall translation quality of GPT-4o is high, with most languages achieving average scores above 4. Notably, performance on Japanese and Thai is comparatively lower, though still above 3.5, suggesting acceptable translation quality for these languages.
To further evaluate the robustness of our framework with respect to translation quality, we conducted a supplementary experiment using the open-sourced LLM Qwen 3-8B [qwen3technicalreport]. Qwen 3-8B achieves a FLORES chrF score of 56, which is lower than GPT-4o’s score of 62 for the same language set. We employed both Qwen 3-8B and GPT-4o to translate our training preference pairs and trained separate MuRater models based on each translation. Each MuRater was then applied to score 10,000 multilingual documents per language. We compared the annotation outputs between the two models using Pearson correlation and Kendall’s Tau.
Table 4: Agreement between MuRater models trained with translations from Qwen 3-8B vs. GPT-4o, measured by Kendall’s Tau and Pearson correlation.
Language ja de es ar id pt th fr vi
Kendall’s Tau 0.8930 0.9005 0.8951 0.8645 0.8903 0.8964 0.8834 0.8924 0.8867
Pearson Corr. 0.9835 0.9854 0.9843 0.9793 0.9822 0.9848 0.9803 0.9831 0.9825
Language it ko ms tl ru tr zh nl
Kendall’s Tau 0.8851 0.8885 0.8427 0.8749 0.8942 0.8822 0.9044 0.8991
Pearson Corr. 0.9815 0.9821 0.9644 0.9755 0.9850 0.9818 0.9861 0.9843
As shown in Table 4, both metrics indicate consistently high agreement across all evaluated languages. These findings suggest that even when relying on a weaker translation model such as Qwen 3-8B, MuRater can still be effectively trained, provided that the relative preference information is preserved. This demonstrates the robustness of our approach to moderate variations in translation quality.
6.6 Pointwise Score
7 Experiment Setup Details
7.1 Dataset
We use the 16 recent snapshots from FineWeb-2 as our raw data before MuRater and other baselines annotation, namely CC-MAIN-2021-39, 2021-43, 2021-49, 2022-05, 2022-21, 2022-27, 2022-33, 2022-40, 2022-49, 2023-06, 2023-14, 2023-23, 2023-40, 2023-50, 2024-10, and 2024-18.
To analyze domain composition, we employ NVIDIA’s multilingual domain classifier [^8] https://huggingface.co/nvidia/multilingual-domain-classifier to annotate the domain distribution of our dataset. Figures 711 illustrate the domain shifts before and after applying the MuRater-based selection. The results show that MuRater systematically prioritizes World Knowledge domains such as People and Society, Health, and Science, which are typically well-structured and rich in informational content—properties particularly beneficial for large language model pretraining. However, the resulting domain distributions vary across languages, primarily reflecting intrinsic differences in the domain composition of their respective source corpora.
We adopt the open-sourced FastText language identification model [joulin2016bag], which supports 176 languages and is widely used in large-scale multilingual data pipelines. For multilingual data selection, we retain the top 10% of highest-scoring documents per language from the raw corpus, preserving the natural data composition of Common Crawl and maintaining a distribution similar to full set of FineWeb-2. The initial token ratio across languages is: ru (14.29%), es (12.83%), ja (12.19%), de (12.19%), zh (9.26%), fr (8.98%), it (7.29%), pt (4.58%), nl (4.53%), vi (3.21%), id (2.88%), ar (2.75%), tr (2.20%), th (1.51%), ko (1.41%), tl (0.04%), ms (0.02%).
We then apply temperature-based sampling with $\tau=3.33$, following standard multilingual pretraining practices [xue2020mt5]. The final token ratios used in training are: ru (8.08%), es (7.78%), ja (7.62%), de (7.62%), zh (7.00%), fr (6.93%), it (6.50%), pt (5.71%), nl (5.70%), vi (5.07%), id (4.86%), ar (4.78%), tr (4.52%), th (4.15%), ko (4.11%), tl (2.75%), ms (2.62%).
@@ 7.1 Dataset
7.2 Baselines
We follow the same annotation procedure for the English datasets of QuRater, AskLLM, DCLM, and FineWeb-Edu as described in Appendix 6. For QuRater-M, we apply the same prompting approach (also detailed in Appendix 6) and instruct GPT-4o to annotate 300,000 multilingual pairs, focusing exclusively on content regardless of the language. We then fine-tune the multilingual QuRater baseline using both English and multilingual data, leveraging the BGE-M3 model [bge-m3] and the identical training hyperparameters outlined in Appendix 6.
7.3 Model Architecture
We utilize a transformer architecture based on the LLaMA-2 model [touvron2023llama], configured to contain approximately 1.2 billion parameters. Models are randomly initialized before pretraining. The detailed information for the model configuration and training hyperparamters is shown in Table 5.
@@ 7.3 Model Architecture
8 Evaluation Benchmarks
All task evaluations are conducted using the lm-evaluation-harness framework [eval-harness]. For English in-context learning tasks, we use the following benchmakrs:
  • ARC-Easy and ARC-Challenge [clark2018think] (25-shot): Multiple-choice science questions from grade school exams, assessing models’ ability to apply scientific knowledge and reasoning.
  • SciQ [welbl-etal-2017-crowdsourcing] (0-shot): Crowdsourced multiple-choice science questions covering physics, chemistry, and biology, designed to evaluate scientific understanding.
  • LogiQA [liu2020logiqa] (0-shot): Logical reasoning questions derived from Chinese civil service exams, testing deductive reasoning capabilities.
  • TriviaQA [joshi2017triviaqa] (5-shot): Reading comprehension dataset with question-answer pairs authored by trivia enthusiasts, accompanied by evidence documents.
  • BoolQ [clark2019boolq] (5-shot): Yes/no questions with associated passages, evaluating models’ ability to answer naturally occurring questions.
@@ 8 Evaluation Benchmarks
  • HellaSwag [zellers2019hellaswag] (10-shot): Sentence completion tasks requiring commonsense inference to select the most plausible continuation.
  • PIQA [bisk2019piqa] (5-shot): Physical commonsense reasoning questions, focusing on everyday tasks and interactions.
  • OpenBookQA [OpenBookQA2018] (10-shot): Multiple-choice questions based on elementary science facts, requiring both factual knowledge and reasoning.
  • WinoGrande [sakaguchi2019winogrande] (5-shot): Pronoun resolution tasks designed to test commonsense reasoning at scale.
Additionally, two World Knowledge tasks are evaluated:
  • Natural Questions (NQ) [kwiatkowski2019natural] (5-shot): Real user questions paired with answers from Wikipedia, assessing open-domain question answering.
  • MMLU [hendrycks2021measuring] (5-shot): A benchmark covering 57 subjects across various domains, measuring multitask language understanding.
For evaluating translated benchmarks, we use the MuBench dataset [han2025mubenchassessmentmultilingualcapabilities] and conduct evaluations across 18 languages present in our training set. In the multilingual setting, we evaluate:
@@
  • StoryCloze [mostafazadeh2016corpus] (0-shot): Narrative understanding task where models choose the correct ending to a four-sentence story.
  • BMLAMA [qi-etal-2023-cross] (0-shot): Multilingual factual knowledge probing dataset, assessing cross-lingual consistency in language models.
  • XCOPA [ponti2020xcopa] (5-shot): Causal commonsense reasoning tasks translated into multiple languages, evaluating cross-lingual inference.
  • XNLI [conneau2018xnli] (5-shot): Cross-lingual natural language inference benchmark, testing entailment and contradiction detection.
  • XWinograd [tikhonov2021s] (5-shot): A multilingual benchmark for evaluating localized knowledge and reasoning abilities of large language models across diverse languages.
  • MultiLoKo [hupkes2025multiloko] (5-shot): A large-scale multilingual evaluation suite designed to assess factual knowledge, reasoning, and question answering across 45 languages, emphasizing both cross-lingual consistency and language-specific understanding.
  • FLORES [goyal2022flores] (5-shot): Multilingual machine translation benchmark, evaluating translation quality across diverse languages.
@@
9 Detailed Results
9.1 English Detailed Results
@@
9.2 Multilingual Detailed Results
Detailed results of different data-selection methods across individual downstream benchmarks. For the 7B experiments, we additionally include the MultiLoKo benchmark [hupkes2025multiloko] to assess cultural and regional knowledge across languages. Since its scores were too low to be meaningful under the 1.2B training setup, we do not report the results here.
Table 7: Performance of different data-selection strategies across downstream tasks when mixing 200B English and 300B multilingual tokens. MuRater(M) denotes scoring multilingual pairs translated into English, while MuRater(E) uses rated English data translated into multilingual pair form. Best results within each setting are shown in bold.
Selection Method MMLU (L) ARC C_ML ARC E_ML FLORES Hella swag_ML MMLU (T) XCOPA XNLI Story Cloze_ML XWino BMLAMA Average 18 Languages Uniform 29.98 28.74 49.03 46.66 44.56 27.83 64.60 42.33 69.28 76.40 48.55 48.00 HPLT-2 29.24 26.94 45.95 42.66 39.87 27.58 59.67 42.14 65.09 71.77 48.38 45.39 FineWeb-2 28.33 27.50 45.97 44.52 42.39 27.75 62.57 41.00 66.77 72.89 41.51 45.56 QuRater-M 30.86 33.89 57.53 46.60 46.77 29.18 62.97 44.39 65.86 71.21 45.84 48.65 MuRater(M) 31.86 34.26 58.21 47.43 46.54 29.28 64.43 42.07 67.08 72.92 50.13 49.47 MuRater(E) 31.96 35.01 58.98 47.27 47.11 29.40 65.73 44.46 67.67 74.13 52.97 50.43 13 Languages Subset FineWeb2-HQ 30.52 30.91 54.37 50.97 46.15 28.66 64.92 41.79 69.48 68.72 43.08 48.14 MuRater(E) 31.96 36.03 61.63 50.11 49.17 29.43 67.44 44.49 68.75 68.83 53.19 51.00
Table 8: Comparison of MuRater and QuRater-M when training a 7B model on 1T tokens with 16.5% multilingual data.
MMLU_L ARC_C_ML ARC_E_ML FLORES HSWAG MMLU_T XCOPA XNLI S. Cloze XWino BMLAMA MultiLoKo Avg.
QuRater-M 35.93 42.36 65.20 55.15 57.29 32.54 69.27 45.13 74.50 82.60 49.61 8.87 51.54
MuRater 36.87 43.38 66.87 55.38 57.76 32.93 71.03 45.06 75.42 83.19 52.82 10.61 52.61
We also display the detailed results of each benchmark and each language below.
@@ 9.2 Multilingual Detailed Results
9.3 Impact of Translation and Data Quality on Multilingual Performance
To further investigate translation and data quality impact on multilingual performance, we analyze two representative language pairs with comparable token ratios but differing translation quality: Japanese vs. Spanish and Thai vs. Korean. Human evaluation results reveal that Japanese and Thai translations receive approximately 0.5 lower average quality scores than their respective counterparts, Spanish and Korean. As detailed in Tables above, this discrepancy is reflected in downstream performance, where Japanese and Thai consistently underperform across most multilingual benchmarks.
Table 20 presents the normalized top-10% selection scores (relative to Chinese). These results show that Japanese and Korean data exhibit notably lower selection scores than Spanish and Thai, aligning with observed translation-quality trends.
Table 20: Normalized top-10% document scores across languages (relative to Chinese).
Language ja de es ar id pt th fr vi
Score 0.743 0.862 0.922 0.877 0.996 0.896 0.922 0.806 0.967
Language ms tl it ko ru tr zh nl
Score 0.817 0.690 0.877 0.843 0.940 0.941 1.000 0.862
These findings suggest that translation quality partially accounts for the observed performance gaps, yet it is not the sole determinant. Additional factors—including the intrinsic quality of the source corpora, language-specific tokenizer compression effects [goldman2024unpacking], and cross-lingual knowledge transfer dynamics [he2024scaling]—likely contribute to the variation in multilingual model performance. Collectively, the results highlight that improving translation fidelity and ensuring balanced corpus quality are both critical for enhancing multilingual LLM training.
10 Case Study
@@
Figure 14: Sampled training examples of French with quality ratings at different score range